OpenStudy (anonymous):

medal+Fanning Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

Directrix (directrix):

This question has to do with the discriminant of a quadratic equation. That becomes a concern when solving a quadratic equation using the Quadratic Formula. Attached for three items for you to study later. The discriminant is called the discriminant because it discriminates among the types of roots a quadratic has. You can predict the type roots based on the discriminant.

OpenStudy (anonymous):

Thank you and i need a numeric answer

Directrix (directrix):

@HelloGoodmorning I get that you need an answer but I thought you might want to know something about how to get the answer for later. If you looked at the attachments, you might know the answer to your first question. Do you know? If so, post that, and then we'll move to the second part. Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring.

Directrix (directrix):

Which one of these three: −b b2 − 4ac 2a

OpenStudy (anonymous):

Loll I'm trying to check it out but my laptop is so slow that i'm unable to

OpenStudy (anonymous):

and what do you mean by which one of these three

Directrix (directrix):

You wrote the question yourself with these three choices: Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a

Directrix (directrix):

These are the three choices: Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a

OpenStudy (anonymous):

Loll I'm trying to check it out but my laptop is so slow that i'm unable to

Directrix (directrix):

This option should read: b² − 4ac

OpenStudy (anonymous):

Ok

OpenStudy (littlebird):

Now just plug in the numbers where the letters are. :)

Directrix (directrix):

2 x2 + 7x + 3 = 0 a = 2, b = 7, and c = 3 b² − 4ac = (7)² - 4 (2)(3) = ? @HelloGoodmorning

OpenStudy (anonymous):

I'm easily confused I have adhd so sorry @Directrix

OpenStudy (anonymous):

Give me a minute ill answer that

Directrix (directrix):

No rush at all. Take your time. Your task is to crank out the value of this expression: (7)² - 4 (2)(3) So multiply 7 times 7. Multiply 4 times 2 times 3. Then, subtract what you got for (4 times 2 times 3) from what you got for (7 times 7).

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

7*7=49 4*2=8 and 8*3 equals =24 so 49-24=25

OpenStudy (anonymous):

So the numerical answer is 25?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Correct.

OpenStudy (anonymous):

Yes! All I have to put in is 25 Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 Numerical Answers Expected! Answer for Blank 1: 25

Directrix (directrix):

Answer for blank 1: b² − 4ac Answer for blank 2: 25

Directrix (directrix):

@HelloGoodmorning

Directrix (directrix):

Wait a minute. How many blanks do you have for filling in. If there is only one and it is numerical, then it is the 25.----> as you said.

OpenStudy (anonymous):

Loll yes there's only one blank

Directrix (directrix):

Put 25 in it and call it good.

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

One question How can you tell when a quadratic equation has two identical, rational solutions? (choices) a. when the radicand is negative b. when b in the quadratic formula is greater than the radicand c. when the radicand equals zero d. when the radicand is not a perfect square

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

I think the answer is c am I correct?

OpenStudy (anonymous):

Your right I'm sorry

Directrix (directrix):

>> I think the answer is c am I correct? Yes.